# Assemblage: Automatic Binary Dataset Construction for Machine Learning

### v1

@@ 4 The Assemblage Datasets
We constructed a dataset of 890K Windows PE binaries as follows: starting from 4 million crawled GitHub repositories dating from 2008 to 2023, we discarded repositories that had fewer than 10 files, were smaller than 50KB in size, lacked a README.md file containing the string “Visual Studio”, or lacked a solution file supporting Microsoft Visual Studio tool chains. We were left with 179,548 repositories, from which Assemblage was then able to build and store binaries using our Microsoft Visual Studio toolchains. Build configurations varied in toolchain version (Visual Studio 2015, 2017, 2019, and 2022), optimization level (0d, 01, 02, 0x), CPU architecture, and Debug/Release mode. In sum, we ended up with 847,856 executable (.exe) files and 43,507 dynamically-linked library (.dll) files, along with PDB files for each binary. In total, we accumulated 2TB worth of data, of which the executables (.exe and .dll files) took 200GB. To conform with ethical data distribution practices, we recorded license information from crawled repositories when available. Table 2 shows the distribution of licenses present among the GitHub repositories in our corpus; consistent with our observation of many small repositories, 158,624 of the 179,548 repositories have no license.
Table 2: Top licenses in GitHub datasets.
License Windows-GitHub Linux-GitHub
MIT 7258 6762
GPL 3078 2845
Apache 1168 859
BSD 313 609
CC 94 50
BSL 22 49
Figure 2: Binary size distribution of Windows on GitHub, vcpkg and Linux on GitHub data.
@@
We measured our dataset along a variety of dimensions to demonstrate its quality and diversity. Figure 2 shows a histogram of sizes of binaries from our (a) Windows binaries from GitHub, (b) vcpkg (c) Linux binaries from GitHub corpuses. Our Windows-GitHub binaries were the smallest overall, with an average size of 143KB. Manual inspection revealed many student projects, basic utilities, and starter codes among our GitHub dataset. However, in absolute terms, our Windows-GitHub dataset still contains a substantial number of large binaries, including 26,549 binaries over 0.5MB.
@@
We also present a breakdown of our Windows-Github dataset by compiler and optimization level (Table 4). The number of binaries at each compiler version and optimization is not equal, due to unevenly distributed computing power among Assemblage worker nodes, changes to the availability of repositories, restricted running time, and the limited availability of some configurations in solution files. To further investigate the varying number of binaries under each build configuration, we recorded the error message encountered during build failures, which we summarize in Table 3. Over the whole dataset, the most common causes of build failure are related to compiler infrastructure, such as invalid solution configurations, Windows SDK errors due to upgrading, invalid or missing project files, and other compilation errors. Although we tried to mitigate the errors caused by missing files by installing a wide range of SDK libraries, complete coverage was infeasible, and thus certain library versions were absent from our worker environment. We also did our best to minimize invalid configuration errors by minimally modifying variables in vcproj files and using devenv from MSVC to upgrade solution files when appropriate.
Table 4: Configuration distribution on three datasets.
OS Source Compiler Opt. Count
Windows PE GitHub MSVC-v140 O1 60126
O2 165866
Ox 100831
MSVC-v141 O1 6113
O2 78203
Od 86841
Ox 7470
MSVC-v142 O1 87497
O2 50055
Od 72613
MSVC-v143 O1 80243
Od 57011
Ox 37302
vcpkg MSVC-v120 O1 952
O2 945
Od 956
Ox 926
MSVC-v142 O1 2160
O2 2142
Od 2187
Ox 2145
MSVC-v143 O1 3081
O2 3078
Od 3074
Ox 3083
Linux ELF GitHub GCC-11.4.0 Od 50338
O1 49860
O3 50995
Oz 22030
Clang-14.0.0 Od 50338
O1 63188
O2 65934
O3 69082
@@
Figure 3: Function distribution of each binary.
We also inspected the function information corresponding to our Windows-Github PE dataset. We found that out of 252M functions, there are 14M unique function names, and 138M unique function hashes. Assemblage is also able to store source code for 17M functions in its SQL database. We also observed significant variation in the number of functions per binary, which we illustrate in Figure 3. Finally, we examined the relationship between a binary’s size and the functions it contains. Counter-intuitively, the number of functions in a binary is not correlated with the size of that binary, as these variables have a Pearson correlation coefficient of 0.22. Furthermore, the size of functions within a binary has a 0.49 correlation coefficient to the binary size, but no correlation to the number of functions in the binary. We unexpectedly conclude that binary size doesn’t have a strong correlation with either the number of functions nor the average function size in a binary. Moreover, the existence of function level information provides insights into the behavior patterns of specific compilers. For example, in 2017 Visual Studio implemented stack-based buffer protection by inserting security cookies, and we can locate more than 744k __security_init_cookie functions with 151k unique hashes. These function level information provided by Assemblage presents valuable data for not only machine learning research, but also a potential of application to a wide range of research areas like security and binary analysis.

### v2

@@ 3.2 Dataset Details and Distribution
We constructed a dataset of 890K Windows PE binaries as follows: starting from 4.6 million crawled GitHub repositories dating from 2008 to 2023, we discarded repositories that had fewer than 10 files, were smaller than 50KB in size, lacked a README.md file containing the string “Visual Studio”, or lacked a solution file supporting Microsoft Visual Studio tool chains. We were left with 179,548 repositories, from which Assemblage was then able to build and store binaries using our Microsoft Visual Studio tool chains. Build configurations varied in toolchain version (Visual Studio 2015, 2017, 2019, and 2022), optimization level (0d, 01, 02, 0x), CPU architecture, and Debug/Release mode. In sum, we ended up with 847,856 executable (.exe) files and 43,507 dynamically-linked library (.dll) files, along with PDB files for each binary. In total, we accumulated 2TB worth of data, of which the executables (.exe and .dll files) took 200GB. To conform with ethical data distribution practices, we recorded license information from crawled repositories when available.
Figure 3: Binary size, Windows PE (GitHub), vcpkg, and Linux ELF (GitHub).
Figure 4: Number of functions in each binary.
We measured our dataset along a variety of dimensions to demonstrate its quality and diversity. Figure 4 shows a histogram of sizes of binaries from our (a) Windows binaries from GitHub, (b) vcpkg and (c) Linux binaries from GitHub corpora.
@@
Our Windows-GitHub binaries were the smallest overall, with an average size of 143KB. Manual inspection revealed many student projects, basic utilities, and starter codes in our GitHub dataset. However, in absolute terms, our Windows-GitHub dataset still contains a substantial number of large binaries, including 26,549 binaries $\geq$ 0.5MB. Note that compiling random Windows projects is non-trivial due to a lack of standardization; more details are provided in Appendix B.
We also inspected the function information corresponding to our Windows-GitHub PE dataset. We found that out of 252M functions, there are 14M unique function names, and 138M unique function hashes. Assemblage was also able to store source code for 17M functions in its SQL database. We also observed significant variation in the number of functions per binary, which we illustrate in Figure 4. Finally, we examined the relationship between a binary’s size and the functions it contains. Counter-intuitively, the number of functions in a binary is not correlated with the size of that binary, as these variables have a Pearson correlation coefficient of 0.22. Furthermore, the size of functions within a binary has a 0.49 correlation coefficient to the binary size, but no correlation to the number of functions in the binary. We unexpectedly conclude that binary size doesn’t have a strong correlation with either the number of functions nor the average function size in a binary. Moreover, the existence of function-level information provides insights into the behavior patterns of specific compilers. For example, in 2017 Visual Studio implemented stack-based buffer protection by inserting security cookies, and we can locate more than 744k __security_init_cookie functions with 151k unique hashes. These function level information provided by Assemblage presents valuable data for not only machine learning research but also a potential application to a wide range of research areas like security and binary analysis.
25k Windows vcpkg Dataset
Upon examining our Windows-GitHub dataset, we observed that only 4.7% of the corpus consisted of library files. To address this deficiency and provide coverage of Windows’ statically-linked and dynamically-linked libraries, we adapted Assemblage to build binaries from vcpkg [48], a dataset of 2,185 actively-developed, well-maintained, and widely-used packages. vcpkg consists of many commonly used and permissively licensed C++ libraries, including graphics and UI-related libraries.
Our vcpkg worker takes advantage of the native commands of vcpkg, and we modified the package manager’s original build script to pass custom variables and flags to its compiler. In this way, we harvested a diverse Windows PE library corpus from vcpkg, comprising 24905 binaries, along with their pdb files, from 2078 different projects. Only 110 of these projects are also present in our Windows-GitHub dataset, demonstrating the added diversity that is present in our vcpkg dataset. Similarly to the Windows-GitHub dataset, vcpkg binaries were built using 4 different versions of the Microsoft Visual Studio toolchain, as well as with 4 optimization levels (-O{d,1,2,3}).
Finally, as illustrated in Fig. 4, the size of the binaries in our vcpkg dataset is larger than the size of those in our Windows-GitHub dataset. The vcpkg binaries were 2124 KB on average, compared to 143KB for Windows-Github binaries, with a small number (roughly 500) of vcpkg binaries $\geq$ 10 MB.
Proof-of-Concept Linux Dataset
Motivated by the scarcity of corpora of benign Windows PE binaries, we have focused our efforts on building such executables. However, to ensure Assemblage’s API was sufficiently expressive, we also added several Linux builders. Our prototype Linux crawler traverses GitHub and searches for Makefiles to identify potential Linux ELF projects. Among 4 million crawled repositories, Assemblage recognized 367,207 repositories that include a Makefile. We created 8 build configurations, using GCC (-O{0,1,2,Z}) and Clang (-O{0,1,2,3}).
In addition to the GitHub crawler, we also built a proof-of-concept crawler and builder that leverages the Arch User Repository (AUR) for source code. Users can set the configuration by providing custom makepkg.conf for each project. While the AUR repositories are well maintained by the community and open-sourced, due to the rate limits and ethical concerns, we did not deploy builders to generate a large scale dataset from AUR. However, our proof-of-concept crawler demonstrates that Assemblage can be configured to work with a variety of package managers.
Assemblage Dataset Distribution Format
We only release the binaries built from repositories that clearly state their license, and a typical Assemblage dataset consists of both (a) an SQLite database of metadata and (b) an archive of binaries (either as a .tar or manually via the Amazon S3 interface). The SQL database provides a logical specification of a collection of binary files, currently .exes and .dlls. The user queries the SQLite database to perform application-specific feature extraction (e.g., identifying pairs of sibling functions for triplet learning), and often feeds the binaries into a machine learning framework such as TensorFlow, PyTorch, or Keras.
Figure 5: Dataset database and usage
The SQLite database consists of four primary tables: binaries, functions, rvas, and lines (illustrated Figure 5). The binaries table retains a reference of each binary’s configuration used during compilation, e.g. platform, build mode, compiler version, and optimization, and assigns each binary a unique identifier for further reference. We also register additional information about each executable including the file name, the location of the project on GitHub, the project’s license, and the size of the binary in this table. Using the binaries table, researchers can curate subsets of the dataset to meet their training needs, e.g. by selecting only binaries of a certain size or with a certain license.
To provide function-level details and minimize the need to distribute PDB files (since, as we saw earlier, PDB files can take up 90+of dataset space), the functions, rvas and lines tables reference the primary key for each binary stored in binaries table to relate functions and binaries. The functions tables offers a high-level overview of functions, presenting a hash of the function’s bytes and the function name, whereas the rvas and lines tables provide lower-level function details. Specifically, the rvas table documents the relation between each function and its relative virtual address when the binary is loaded into memory. On the other hand, the lines table illustrates the mapping of each source code line to the precise byte address located in the compiled binary. The dataset also provides the relation between binaries and their corresponding PDB files in a pdbs table, in case researchers want to make use of those files for further binary analysis.
Figure 5 shows the information that is available about a single binary within a Assemblage dataset. As an example, we query the database for a binary called butler.exe, which has the unique identifier 184456.
In the binaries table, we get basic facts about the executable, including the compiler version (Visual Studio 2017), build mode (Release), optimization flag (O2), and associated license (MIT). In the functions table, we can see that this binary has 568 functions; in particular, we focus on one function called tinyxml2::StrPair::SetStr, uniquely identified by the tag 51629995. We use this unique identifier to query the lines and rvas tables for further information about this function. In the lines table, we can see the source code corresponding to this function, such as the line if (*(pu + 0) == TIXML_UTF_LEAD_0. From the rvas table, we learn that this function will be loaded into the relative addresses from Ox6460 to Ox64AA. Thus, this example demonstrates how all the information needed to do binary analysis with corpuses of Assemblage executables is available in the accompanying Assemblage SQLite database.
Ethical Concerns It is natural to be concerned about the presence of malware in any corpus of binary executables. Assemblage builds sources from GitHub, which prohibits malware and allows users to report suspected malware to its community discussion board. Currently, Assemblage does not independently check for malware; we are looking to add malware scanning, following the model pioneered by EMBER [4], in the next iteration of the tool. In addition, the public GitHub repository links are included in our metadata (we do not provide author-level metadata, but plan to distribute this in the future in normalized form), which links the binary data to the author’s GitHub profile. Our public datasets include only those repositories for which permissive licenses may be obtained; we facilitate archiving corpora of unlicensed repositories by distributing “recipes.”
4 Benchmarks
In this section, we demonstrate the utility of Assemblage for a variety of machine learning tasks. Three widely-regarded machine learning approaches to binary analysis problems are evaluated under their original protocols, baring the minimum possible changes to support/read Windows PE files over the original Linux files: a LightGBM model for compiler provenance identification [21], a graph neural network (GNN) for binary function similarity detection [42], and a BERT-based transformer network for understanding binary files, fine-tuned on the task of binary function similarity detection [66]. These experiments broadly show that the strong results of prior literature do not yet generalize to Windows and more diverse compilation settings, and that Assemblage provides the data to further explore these topics (amongst others).
Experiments were run on an Ubuntu 22.04 server equipped with an AMD EPYC 7713P@1.9Ghz 64-Core Processor and 512GB of RAM.
4.1 Compiler Provenance
Figure 6: PassTell’s Confusion Matrix on the dataset mixing the author’s training data (top left) with Assemblage Windows PE binaries (bottom right).
Compiler provenance inference is the task of determining the compiler configuration used to build a binary, including compiler version, build flags, and compiler passes. We trained a state-of-the-art provenance learning model PassTell [21] on a combination of data, mixing the authors’ original Linux dataset with Windows PE binaries obtained from Assemblage. We first used Assemblage to build a dataset of Windows binaries that complemented the Linux binaries studied in the original PassTell paper. We attempted to build each of the 15 source projects selected by Du et al. [21], but due to differences between the Linux and Windows platforms and their respective compiler infrastructures, some projects could not be rebuilt. In total, we collected 127 binaries, which we proceeded to preprocess via the same protocol as the original paper. Though the PassTell authors used objdump, a linear disassembler, to extract function bytes, we found higher success rates with IDA Pro, augmented with the PDB files collected by Assemblage, to accurately handle the significant number (5%) of discontiguous functions present in Windows binaries. Following disassembly, we masked operands and literals in the same manner as Du et al. [21], and balanced the training data by compiler toolchain, version, and optimization level.
We evaluate PassTell twice, once using the aforementioned mix of original Linux and new Windows PEs and a second time using only Windows PEs. The confusion matrix in Figure 7 shows that the PassTell approach over-fits the original Linux models (top left diagonal), but has greater difficulty with the Windows PE files (bottom right). To determine if this is a capacity issue, we re-evaluate on only the Windows PEs to reduce the number of classes and make the ML problem easier, but Figure 7 shows that PassTell still struggles on Windows PEs with $\approx 58\%$ accuracy. The observation of PassTell’s performance on Windows data indicates the limits of training models on small binary datasets lacking diversity. While the approach put forth by Du et al. [21] seemed promising when applied to a small Linux dataset, extending the work to Windows data from Assemblage revealed serious shortcomings. This underscores the utility of Assemblage as a tool for machine learning researchers in the binary analysis space.
4.2 Binary Function Similarity
Table 2: AUC scores evaluating binary function similarity model generalization between the Linux and Windows domains. Column names indicate train/test platform.
Task
Windows/
Linux
Linux/
Windows
Linux/
Linux
Windows/
Windows
arch 0.50 0.97
bit 0.70 0.99
comp 0.63 0.80
opt 0.72 0.62 0.88 0.89
ver 0.82 0.64 0.98 0.84
XA 0.48 0.86
XC 0.63 0.86
XC+XB 0.61 0.87
XM 0.54 0.61 0.87 0.86
Table 3: Reproducing jTrans on Assemblage Windows and Linux datasets, measuring in MRR (higher is better).
Training Epochs 1 20 50
Fine-tune & Evaluate on Windows 0.17 0.52 0.52
Fine-tune & Evaluate on Linux 0.83 0.83 0.83
Fine-tune & Evaluate on Linux (BinaryCorp-26M) 0.82 - 0.98
Next, we used Assemblage to evaluate how well a graph neural network (GNN) trained on Linux data generalizes to Windows binaries.
We conducted these experiments on the GNN developed by Li et al. [42] for binary function similarity detection, as implemented and open-sourced by Marcelli et al. [47]. [^2] Code from https://github.com/Cisco-Talos/binary_function_similarity/.
We made only one change to their data processing pipeline — instead of removing singleton functions, we allowed them to stay in the dataset and be part of negative pairs, as they can still provide a training signal to the model this way and improved performance in a small-scale test.
Using Assemblage and Marcelli et al. [47]’s protocol, we end up with 1.78 million functions for triplet learning. Negative pairs consist of two different functions, whereas positive pairs consist of functions originating from the same source code, but differing in one or more of the following: architecture, bitness, compiler, compiler version, and optimization. Many tasks/pairs that make sense in Linux (e.g., ARM vs. PowerPC and endianness changes) are not widely applicable in Windows PE code bases, and so can not be evaluated in this context. For this reason, Table 2 shows Linux evaluation results on the left side, the original results in Linux/Linux column, and Windows evaluation results on the right for the viable tests. Irrespective, we see that despite the cross-platform/functional goals of Marcelli et al. [47], the dichotomy in Windows PEs and more diverse datasets results in dramatically lower performance. The GNN model trained on Assemblage Windows data shows that, despite using the same code and sophisticated disassembly tools, the gap in performance is non-trivial and an open problem for future study.
A second experiment using the BERT [20] based jTrans [66] is considered. jTrans uses masked language modeling and a novel task, jump target prediction, in its pre-training phase. Despite the jump target prediction code released by Wang et al. [66] requiring Linux specific tooling, all information needed to run their fine-tuning was already in Assemblage’s SQLite database. Otherwise, we precisely followed the script released by the jTrans authors to create a training and validation set.
We randomly selected 72,000 Windows PE binaries and 23,000 Linux ELF binaries from Assemblage’s dataset, processed them, and separated them into training and validation sets for Windows and Linux, respectively. For both the Windows and Linux datasets, using the hyperparameter settings recommended by the jTrans authors, we ran fine-tuning for 50 epochs, at which point the Mean Reciprocal Rank (MRR) over the training set became stable. Table 3 shows a notable difference between the reported MRR scores on Linux (0.83) and Windows (0.52) data, even with fine-tuning. Because the authors have not released their code for pre-training, we could not conduct any further investigation into the impact of the pre-training step. However, it is significant that jTrans and jTrans-zero (jTrans without fine-tuning) both perform worse on the larger, more diverse Assemblage data than they do on their original training corpuses. This demonstrates the value that Assemblage brings to researchers, who can leverage its datasets to conduct experiments on more challenging and more realistic data.
4.3 Binary Function Identification
Table 4: F-1 scores of XDA and IDA on Windows PE binary function boundary identification
Dataset XDA IDA
IDA
(w/PDBs)
WinPE 0.75 0.47 0.79
vcpkg 0.81 0.86 0.86
[…]
5 Conclusion
At first blush, binary analysis fits into a common application pattern of deep learning: an inversion problem where the compiler’s process of converting source code into binary files is the easy generating pass, and machine learning methods are used to learn the inversion from binary back to source code. ML research in binary analysis has thus exploded, but has often not translated to practice due to a gap between “in-the-lab” code and “in-the-wild” binaries. Assemblage helps to mitigate this gap and provides an extensible system that can generate the data needed to push this needed research into the Windows domain, which has the greatest need and the least data/study. This has been shown for three subdomains of binary analysis of compiler provenance, function similarity, and function identification. In each case, we see Assemblage makes it easy to set up experiments following existing methods, obtain new results, identify areas for improvement, and ultimately stable Assemblage’s general purpose utility for binary analysis research.
Acknowledgments
We would like to acknowledge the feedback of the anonymous NeurIPS dataset track reviewers. Additional computing infrastructure (GPUs and servers) was provided by Syracuse University.
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes]
    2. (b)
      Did you describe the limitations of your work? [Yes] Discussed in Ethical Concerns in Section 3.2
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [Yes]
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A]
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A]
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] Included in supplemental materials and appendix.
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes]
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [N/A]
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] We provide AWS deployment detail in supplemental materials.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes]
    2. (b)
      Did you mention the license of the assets? [Yes] Section 1 and Section 3.2
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [Yes]
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes] We only release the binaries built from repository with open source license, stated in Section 3.2
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [Yes] We mention in our paper that repositories may include malware and we are reproducing GitHub, we do not have fine-grained author data but are working on a normalized version of it for the future
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A]
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A]
Appendix A Technical documentation and dataset access
The homepage of Assemblage is https://assemblage-dataset.net/, documentation for Assemblage can be accessed at the link https://assemblagedocs.readthedocs.io, and the binary dataset built from repository with license is hosted on Hugging Face and accessible by everyone, all the links are listed in the documentation.
We also release the codes we used to build the dataset on GitHub, link at https://github.com/Assemblage-Dataset/Assemblage. Anyone can obtain the copy of our source code and replicate our dataset with our recipes. We also provide specific instructions on deployment.
Appendix B Note on the difficulty of compiling random projects
We also present a breakdown of our Windows-GitHub dataset by compiler and optimization level (Table 5). The number of binaries for each compiler version and optimization is not equal due to unevenly distributed computing power among Assemblage worker nodes, changes to the availability of repositories, restricted running time, and the limited availability of some configurations in solution files.
Table 5: The distribution of configurations (opt: optimization level). The number of files at each optimization level is realatively even, but outliers exist for a variety of reasons. Projects do not compile effectively under different optimization settings without manual intervention, cause time-out issues in taking too long to compile, or a variety of other factors. How to automatically build more projects is an open question for future work.
Compiler Opt Count
MSVC-v140 O1 64,598
O2 177,531
Ox 107,635
MSVC-v141 O1 6,503
O2 84,237
Od 93,155
Ox 7,547
MSVC-v142 O1 91,501
O2 53,062
Od 76,139
MSVC-v143 O1 85,171
Od 59,582
Ox 40,004
Compiler Opt Count
MSVC-v141 O1 2,864
O2 2,864
Od 2,624
Ox 2,855
MSVC-v142 O1 8,053
O2 8,077
Od 6,230
Ox 7,888
MSVC-v143 O1 3548
O2 3548
Od 3,519
Ox 3,282
Compiler Opt Count
GCC-11.4.0 O0 66,919
O1 56,919
O3 61,907
Oz 27,952
Clang-14.0.0 O0 74,947
O1 75,973
O2 81,093
O3 83,963
To further investigate the varying number of binaries under each build configuration, we recorded the error message encountered during build failures, which we summarize in Table 6. The most common causes of build failure are related to compiler infrastructure, such as invalid solution configurations, Windows SDK errors due to upgrading, invalid or missing project files, and other compilation errors. Although we tried to mitigate the errors caused by missing files by installing a wide range of SDK libraries, complete coverage was infeasible, and thus, certain library versions were absent from our worker environment. We also did our best to minimize invalid configuration errors by minimally modifying variables in vcxproj files and using devenv from MSVC to upgrade solution files when appropriate.